OpenStudy (anonymous):

If (x+2)/2 = y/5, then which of the following must be true? x/2 = (y-2)/5 x/2 = (y-5)/5 (x+2)/5 = y/25 HELP PLEASE!

OpenStudy (anonymous):

If (x+2)/2 = y/5 then x = 2y/5 -2 = 2(y-5)/5. Hence x/2 = (y-5)/5 is true.

OpenStudy (anonymous):

Thank you